# LLM-assisted Explicit and Implicit Multi-interest Learning Framework for Sequential Recommendation

### v1

@@ 3.1. Problem Definition
In this paper, we define the set of all users as $\mathcal{U}$ and the set of all items as =$\mathcal{I}=\{i_{1},i_{2},\cdots,i_{N}\}$, Where N is the number of unique items in the dataset. For each historical interaction sequence $s_{u}$ between a user and an item set, we represent it as $S_{u}=[i_{1}^{u},i_{2}^{u},\cdots,i_{n}^{u}]$. Here $n$ is the maximum length of the sequence, and the sequence is arranged in order according to the time of user interaction. Therefore, the task of the multi-interest SR model is to recall a subset of items from item set $\mathcal{I}$ that the user is likely to interact with next based on the user’s historical behavior records. Specifically, given a user’s interaction sequence $S_{u,t}=[i_{1}^{u},i_{2}^{u},\cdots,i_{t}^{u}]$ at the previous $t$ steps as input, predict the item $i_{t+1}^{u}$ that the user may click at the next step.
@@ 3.2.1. Dividing Interest Groups
Although each user’s interests are highly individualized, in practical applications, those users who exhibit similar behavioral patterns are likely to share similar interests. Second, considering the relationship between data volume and LLM inference time, to effectively reduce the time required for the LLM inference process, we start from the perspective of data optimization by selecting representative data to ensure the quality of the model inference and significantly reduce the computational resource requirements.
Based on the natural similarity of text structure, we cluster the sequence data in text form to divide users into different interest groups. We chose the Affinity Propagation (AP) algorithm (Frey and Dueck, 2005) as our clustering tool during this process. The uniqueness of this algorithm is that it determines the relationship between data points by exchanging “responsibility” and “availability” messages between data points until a stable state is reached to form clusters. This mechanism enables Affinity Propagation to dynamically determine the optimal number of clusters without pre-setting, thereby more accurately capturing the different interest patterns of users. The specific algorithm flow is shown in Algorithm 1.
(1) $E_{t}=\text{BERT}(T),$
where $T=[t_{1},t_{2},\cdots,t_{n}]$ represents the text form of the user’s historical click sequence, BERT is a pre-trained language model.
(2) $C,C_{\text{c}}=\text{AP}(E_{t},p),$
where $C=\{c_{1},c_{2},\cdots,c_{K}\}$ represents the category of clustering and $K$ is the number of clusters, and $p$ is the hyperparameter in the AP algorithm, which can control the scale of clustering. $C_{\text{c}}=\{c_{\text{c}}^{1},c_{\text{c}}^{2},\cdots,c_{\text{c}}^{1},c_{\text{c}}^{K}\}$is the central sample set for each cluster, we use these central samples as typical samples for LLM inference. Note that for clusters where the center sample cannot be found, we select the closest-spaced sample as the typical sample by calculating the distance from the center value for all samples within the cluster.
@@ 3.2.2. Constructing Typical Prompts
Algorithm 1 Affinity Propagation Clustering Algorithm.
0:  $\mathbf{E_{t}}\in\mathbb{R}^{n\times d}$, $p$
0:  $\mathbf{C}$, $\mathbf{C_{c}}$
1:  Initialize responsibility matrix $\mathbf{R}\leftarrow\mathbf{0}_{n\times n}$
2:  Initialize availability matrix $\mathbf{A}\leftarrow\mathbf{0}_{n\times n}$
3:  Compute similarity matrix:
4:  $s(i,k)=(-\|E_{t}[i]-E_{t}[k]\|^{2}+p)$ for $i,k$ in range($n$)
5:  while not converged do
6:     Update the responsibility matrix $\mathbf{R}$ using the current availability matrix $\mathbf{A}$:
7:     $r(i,k)\leftarrow s(i,k)-\max_{k^{\prime}\neq k}\{a(i,k^{\prime})+s(i,k^{\prime})\}$
8:     Update the availability matrix $\mathbf{A}$ using the updated responsibility matrix $\mathbf{R}$:
9:     $a(k,k)\leftarrow\sum_{i^{\prime}\neq k}\max\{0,r(i^{\prime},k)\}$
10:     Determine exemplars by finding column $k$ that maximizes $a(i,k)+r(i,k)$ for $i$ in range($n$)
11:  end while
12:  Assign each data point to its nearest exemplar
13:  return  $\mathbf{C}$, $\mathbf{C_{c}}$
@@ 3.2.3. Inference Semantic Interests
where $P_{T_{c}}=\{p_{t_{c}}^{1},p_{t_{c}}^{2},\cdots,p_{t_{c}}^{K}\}$ represents the prompts corresponding to the typical samples. The user’s semantic interest of each sample in $T_{\text{r}}$ is composed of multiple sub-interests $t_{\text{r}}^{j}$.
@@
At the same time, to distinguish each semantic interest and ensure the consistency of text representation, we use the same BERT pre-trained model as in the previous section 3.2.1 to encode each sub-interest separately, and then concatenate them to form a semantic interest representation of a typical sample.
(4) $h_{\text{ex}}=\text{Concat}(\text{Bert}(t_{\text{r}^{u}}^{1}),\text{Bert}(t_{\text{r}^{u}}^{2}),\cdots,\text{Bert}(t_{\text{r}^{u}}^{m})),$
@@
$t_{\text{r}}^{m}\in T_{\text{r}}$ $h_{\text{ex}}\in\mathbb{R}^{m\times d_{t}}$ is the user’s explicit semantic interest representation obtained by LLM inference, $m$ represents the number of interests in LLM inference, and $d_{t}$ is the text embedding dimension. For each user’s semantic interest, we need to use the dictionary stored by the previous clustering algorithm in the form of “$\text{user}_{id}$-($C_{\text{user}_{id}}$, $C_{c})$” to locate the corresponding typical sample semantic interest and use it as the interest representation of the user.
@@ 3.3.1. Multi-interest SR Model
To more effectively mine potential interests from user behavior data represented in the form of IDs, the EIMF framework integrates the interfaces of traditional multi-interest recommendation models. We denote the set of all user ID sequences as $S=[s_{1},s_{2},\cdots,s_{n}]$,
(5) $H_{\text{im}}=\text{Multi-Interest SR Model}(S),$
where $h_{im}\in\mathbb{R}^{n_{i},d}$ is the learned user implicit behavior interest, $n_{i}$ represents the number of interests learned from the sequence and $d$ is the embedding dimension.
@@ 3.3.2. Target-aware Attention Layer
The multi-interest SR model allows us to capture multiple implicit behavioral interests of a user, each of which reflects a different aspect of the user’s behavioral interests. Given that the ultimate goal of the SR model is to predict the user’s next click item, a target-aware attention mechanism is specifically designed. This mechanism uses target labels to screen out the most relevant potential behavioral interests during the training process. Specifically, in the target-aware attention layer, the query is the target label ($e_{i}^{T}$), and $k,v=h_{\text{im}}$.
(6) $\vec{h_{\text{im}}}=T_{a}\textit{-}\text{Attention}(q,k,v),$
@@ 3.4.1. Semantic Prediction Task
Consistent with the design concept of the attention layer in the implicit behavior interest module in the previous section, we designed a two-layer attention mechanism for the explicit semantic interests derived from LLM inference. The self-attention layer aims to learn the association preferences between different semantic interests based on the semantic interests themselves; the target-aware attention layer selects the semantic interests most relevant to the semantic prediction task by injecting the text labels of the target products.
(7) $h_{t}=\text{Self}\textit{-}\text{Attention}(q,k,v),$
where $q,k,v=h_{\text{ex}}$.
(8) $\vec{h_{\text{ex}}}=T_{a}\textit{-}\text{Attention}(q,k,v),$
where $q=e_{t}^{T}$, $k,v=h_{t}$, $e_{t}$ represents the text embedding of the target item name. Then, we calculate the score of user semantic interest embedding and item text embedding $\hat{y_{t}^{k}}=\text{softmax}(\vec{h_{\text{ex}}}^{T}e_{t}^{k}),$ and use cross entropy as the loss function for the semantic prediction task.
(9) $\mathcal{L}_{\text{S}}=\sum_{k=1}^{n}y_{t}^{k}log(\hat{y_{t}^{k}}),$
3.4.2. Modal Alignment Task
To ensure that the embeddings of different modalities can be in a unified embedding space, we specially designed a Modal alignment task. Specifically, the task achieves alignment between behavioral interest representation and semantic interest representation as well as alignment between item ID labels and corresponding text labels through contrastive learning and cosine similarity.
(10) $\text{CL}(e_{a},e_{b})=-\frac{1}{N}\sum_{k=1}^{N}log\left(\frac{\text{exp}(\text{Sim}(e_{a}^{k},e_{b}^{k})/\tau)}{\sum_{j=1}^{N}\text{exp}(\text{Sim}(e_{a}^{k},e_{b}^{j})/\tau)}\right),$
(11) $\text{Cos}(e_{a},e_{b})=\frac{1}{N}\sum_{k=1}^{N}\left(1-\text{Sim}(e_{a}^{k},e_{b}^{k})\right),$
where $\text{Sim}(a,b)=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$is the calculation method of cosine similarity, the loss function for the final modal alignment task we denote as follows,
(12) $\mathcal{L}_{\text{A}}=\alpha(\text{CL}(h_{\text{ex}},h_{\text{im}})+\text{CL}(e_{\text{t}}^{T},e_{\text{i}}^{T}))+\beta(\text{Cos}(h_{\text{ex}},h_{\text{im}})+\text{Cos}(e_{\text{t}}^{T},e_{\text{i}}^{T})),$
where $\alpha$ and $\beta$ are hyperparameters, $\alpha+\beta=0.5$.

### v3

@@ 3.1. Problem Definition
In this paper, we define the set of all users as $\mathcal{U}$ and the set of all items as =$\mathcal{I}=\{i_{1},i_{2},\cdots,i_{N}\}$, Where N is the number of unique items in the dataset. For each historical interaction sequence $s_{u}$ between a user and an item set, we represent it as $S_{u}=[i_{1}^{u},i_{2}^{u},\cdots,i_{L}^{u}]$. Here $L$ is the maximum length of the sequence, and the sequence is arranged in order according to the time of user interaction. Therefore, the task of the multi-interest SR model is to recall a subset of items from item set $\mathcal{I}$ that the user is likely to interact with next based on the user’s historical behavior records. Specifically, given a user’s interaction sequence $S_{u,k}=[i_{1}^{u},i_{2}^{u},\cdots,i_{k}^{u}]$ at the previous $k$ steps as input, predict the item $i_{k+1}^{u}$ that the user may click at the next step. Table 1 shows the symbols and their definitions used in the paper.
3.2. Overview of EIMF
Figure 1 shows the overall architecture of EIMF. From the figure, we can see that the framework is mainly divided into two modules: IBIM and ESIM. In the training stage, EIMF takes the behavioral data in the form of ID and the sequence of user-clicked item names in the form of text as input. The multi-interest SR model in the IBIM module is responsible for learning the implicit interest representation of users in the behavioral data and guides the training process by introducing a label-based target-aware attention layer. In the ESIM module, the AP algorithm first divides users into different clusters based on text similarity, and the cluster centers will be selected as typical samples to construct the corresponding prompt and send to the LLM for explicit interest inference. For the explicit interests obtained by inference, the self-attention layer first learns the association preference between semantic interests and then uses the target-aware attention layer to select the semantic interests most relevant to the semantic prediction task by injecting the text label of the target item. In addition to the recommendation task, EIMF additionally adds modality alignment tasks and semantic prediction tasks to enhance the user behavior interest representation using semantic signals. In the service stage, accurate recommendations can be made based solely on user behavior data.
@@ 3.3.1. Dividing Interest Groups
Directly using LLMs to infer users’ multiple interests from lengthy interaction sequences can be time-consuming. Inspired by the observation that users with similar behavioral patterns often share similar interests, we optimize the input data by selecting representative samples through a clustering approach. This strategy helps maintain inference quality while significantly improving inference efficiency.
Specifically, based on the natural similarity of text structure, we cluster the sequence data in text form to divide users into different interest groups. We chose the Affinity Propagation (AP) algorithm (Frey and Dueck, 2005) as our clustering tool during this process. The uniqueness of this algorithm is that it determines the relationship between data points by exchanging “responsibility” and “availability” messages between data points until a stable state is reached to form clusters. This mechanism enables Affinity Propagation to dynamically determine the optimal number of clusters without pre-setting, thereby more accurately capturing the different interest patterns of users.
(1) $\mathbf{T}=PLM(t),$
where $t=[t_{1},t_{2},\cdots,t_{L}]$ represents the text form of the user’s historical click sequence, $PLM$ is a pre-trained language model. In this paper, we instantiate it using BERT.
(2) $C,C_{\text{c}}=AP(\mathbf{T},p),$
where $C=\{c_{1},c_{2},\cdots,c_{G}\}$ represents the category of clustering and $G$ is the number of clusters, and $p$ is the hyperparameter in the AP algorithm, which can control the scale of clustering. $C_{\text{c}}=\{c_{\text{c}}^{1},c_{\text{c}}^{2},\cdots,c_{\text{c}}^{G}\}$is the representative sample set for each user cluster. We use the advanced LLM to infer semantic interests for these representative samples. After that, for any users, we only need to identify their corresponding representative samples and assign the semantic interests to them, which significantly reduces the inference costs caused by incorporating LLMs. Note that for clusters where the center sample cannot be found, we select the closest-spaced sample as the typical sample by calculating the distance from the center value for all samples within the cluster.
@@ 3.3.3. Inference Semantic Interests
where $T_{\text{r}}=[t_{r}^{1},t_{r}^{2},\cdots,t_{r}^{M_{\text{ex}}}]$ represents the semantic interest of the typical sample text form, which consists of $M_{\text{ex}}$ sub-interests, and $P_{T_{c}}$ represents its corresponding prompt.
@@
At the same time, to distinguish each semantic interest and ensure the consistency of text representation, we use the same pre-trained model as in the previous section 3.3.1 to encode each sub-interest separately, and then concatenate them to form a semantic interest representation of a typical sample.
(4) $\mathbf{H}_{\text{ex}}=\text{Concat}(PLM(t_{\text{r}}^{1}),PLM(t_{\text{r}}^{2}),\cdots,PLM(t_{\text{r}}^{M_{\text{ex}}})),$
@@
where $\mathbf{H}_{\text{ex}}\in\mathbb{R}^{M_{\text{ex}}\times d_{t}}$ is the user’s explicit semantic interest representation obtained by LLM inference, $M_{\text{ex}}$ represents the number of interests in LLM inference, and $d_{t}$ is the text embedding dimension. Note that LLM only infers the semantic interests of typical samples. For each user, it is necessary to find the clustered group and use the semantic interests of the typical samples in the group as their own semantic interests.
Algorithm 1 ESIM: Explicit Semantic Interest Module
1:Input: Text sequences ${T_{u}|u\in\mathcal{U}}$, hyperparameter $p$ for AP clustering
2:Output: Explicit semantic interest representations $\mathbf{H}_{\text{ex}}^{\prime}$ for all users
3:Step 1: Text Encoding
4:Encode text sequences: $\mathbf{T}=PLM(T)$, where $T=[t_{1},t_{2},\dots,t_{L}]$triangleright Section 3.3.1
5:Step 2: Affinity Propagation Clustering
6:Initialize similarity matrix $S$:
7:for each pair $(i,j)$ in $\mathbf{T}$ do
8:     $S(i,j)=-\left|\mathbf{T}(i)-\mathbf{T}(j)\right|_{2}^{2}$triangleright Negative Euclidean distance
9:end for
10:Set preference: $S(i,i)=p$ for all $i$triangleright Control number of clusters
11:Initialize responsibility $R(i,j)=0$ and availability $A(i,j)=0$ for all $i,j$
12:repeat
13:     Update Responsibility:
14:     for each $i,j$ do
15:         $R(i,j)=S(i,j)-\max_{k\neq j}\left\{A(i,k)+S(i,k)\right\}$triangleright Responsibility message
16:     end for
17:     Update Availability:
18:     for each $i,j$ do
19:         $A(i,j)=\min\left\{0,R(j,j)+\sum_{k\neq i,k\neq j}\max\{0,R(k,j)\}\right\}$ for $i\neq j$
20:         $A(j,j)=\sum_{k\neq j}\max{0,R(k,j)}$triangleright Self-availability
21:     end for
22:until convergence or max iterations reached
23:Identify exemplars: $C_{c}={c_{c}^{1},c_{c}^{2},\dots,c_{c}^{G}}$ where $R(i,i)+A(i,i)>0$
24:Assign clusters: $C={c_{1},c_{2},\dots,c_{G}}$ based on nearest exemplar $\triangleright$ Section 3.3.1
25:Step 3: Constructing Typical Prompts
26:for each exemplar $c\in C_{c}$ do
27:     Construct prompt $P_{T_{c}}$:
28:”The user’s historical click sequence is as follows: [item list]; please infer the user’s interest preference and output it in JSON format.” $\triangleright$ Section 3.3.2
29:end for
30:Step 4: Inferring Semantic Interests
31:for each exemplar $c\in C_{c}$ do
32:     Infer semantic interests: $T_{r}=LLM(P_{T_{c}})$, where $T_{r}=[t_{r}^{1},t_{r}^{2},\dots,t_{r}^{M_{\text{ex}}}]$triangleright Section 3.3.3
33:     Encode sub-interests: $\mathbf{H}_{\text{ex}}=\text{Concat}(PLM(t_{r}^{1}),PLM(t_{r}^{2}),\dots,PLM(t_{r}^{M_{\text{ex}}}))$
34:     Compute association preferences: $\mathbf{H}_{\text{ex}}^{s}=\text{Self-Attention}(\mathbf{H}_{\text{ex}})$triangleright Section 3.5.1
35:     Target-aware attention: $\mathbf{H}_{\text{ex}}^{\prime}=T_{a}\text{-Attention}(\mathbf{t}_{\text{tar}},\mathbf{H}_{\text{ex}}^{s},\mathbf{H}_{\text{ex}}^{s})$
36:end for
37:Return: Returns the explicit interest representation of the typical sample in each cluster
Algorithm 2 IBIM: Implicit Behavioral Interest Module
1:Input: Historical interaction sequences $\{S_{u}|u\in\mathcal{U}\}$, target item IDs $\{i_{k+1}^{u}\}$, item embeddings $\{\mathbf{e}_{k}|i_{k}\in\mathcal{I}\}$
2:Output: Implicit behavioral interest representations $\mathbf{H}_{\text{im}}^{\prime}$ for recommendation
3:Step 1: Multi-interest Sequential Recommendation Model
4:Learn implicit interests: $\mathbf{H}_{\text{im}}=MI(S_{u})$, where $S_{u}=[i_{1}^{u},i_{2}^{u},\dots,i_{L}^{u}]$triangleright Section 3.4.1
5:Step 2: Target-aware Attention Mechanism
6:Set query as target item embedding: $q=\mathbf{e}_{\text{Tar}}$, where $\mathbf{e}_{\text{Tar}}$ is the embedding of $i_{k+1}^{u}$triangleright Section 3.4.2
7:Set key and value: $k,v=\mathbf{H}_{\text{im}}$
8:Compute attention weights:
9:$a_{ij}=\frac{\exp(q^{T}k_{j}/\sqrt{d})}{\sum_{m=1}^{M_{\text{im}}}\exp(q^{T}k_{m}/\sqrt{d})}$, where $k_{j}$ is the $j$-th interest vector in $\mathbf{H}_{\text{im}}$
10:Refine representation: $\mathbf{H}_{\text{im}}^{\prime}=\sum_{j=1}^{M_{\text{im}}}a_{ij}v_{j}$, where $v_{j}$ is the $j$-th value vector in $\mathbf{H}_{\text{im}}$
11:Step 3: Recommendation Score Calculation
12:Compute recommendation scores: $\hat{y_{k}}^{\text{t}}=\text{softmax}(\mathbf{H}_{\text{im}}^{\prime}\mathbf{e}_{k})$, for each candidate item $i_{k}\in\mathcal{I}$
13:Return: $\mathbf{H}_{\text{im}}^{\prime}$ for all users
@@ 3.4.1. Multi-interest SR Model
To more effectively mine potential interests from user behavior data represented in the form of IDs, the EIMF framework integrates the interfaces of traditional multi-interest recommendation models. Here, the user ID sequence is $S_{u}=[i_{1}^{u},i_{2}^{u},\cdots,i_{L}^{u}]$,
(5) $\mathbf{H}_{\text{im}}=MI(S_{u}),$
where $\mathbf{H}_{\text{im}}\in\mathbb{R}^{M_{\text{im}}\times d}$ is the learned user implicit behavior interest, $M_{\text{im}}$ represents the number of interests learned from the sequence and $d$ is the embedding dimension. The multi-interest SR model ($MI$) is a sequential recommendation model interface for modeling behavior data.
@@ 3.4.2. Target-aware Attention Layer
The multi-interest SR model allows us to capture multiple implicit behavioral interests of a user, each of which reflects a different aspect of the user’s behavioral interests. Given that the ultimate goal of the SR model is to predict the user’s next click item, a target-aware attention mechanism is specifically designed. This mechanism uses target labels to screen out the most relevant potential behavioral interests during the training process. Specifically, in the target-aware attention layer, the query is the target label ($\mathbf{e}_{\text{Tar}}$) in the form of ID, and $k,v=\mathbf{H}_{\text{im}}$.
(6) $\mathbf{H}_{\text{im}}^{{}^{\prime}}=T_{a}\textit{-}\text{Attention}(q,k,v),$
@@ 3.5.1. Semantic Prediction Task
Consistent with the design concept of the attention layer in the implicit behavior interest module in the previous section, we designed a two-layer attention mechanism for the explicit semantic interests derived from LLM inference. The self-attention layer aims to learn the association preferences between different semantic interests based on the semantic interests themselves; the target-aware attention layer selects the semantic interests most relevant to the semantic prediction task by injecting the text labels of the target items.
(7) $\mathbf{H}_{\text{ex}}^{s}=\text{Self}\textit{-}\text{Attention}(q,k,v),$
where $q,k,v=\mathbf{H}_{\text{ex}}$. After the self-attention layer, we add consideration of the text embedding of the target item through the target attention layer:
(8) $\mathbf{H}_{\text{ex}}^{{}^{\prime}}=T_{a}\textit{-}\text{Attention}(q,k,v),$
where $q=\mathbf{t}_{\text{tar}}\in\mathbb{R}^{d_{\text{t}}}$ represents the text embedding of the target item name, and $k,v=\mathbf{H}_{\text{ex}}^{s}\in\mathbb{R}^{M_{\text{ex}}\times d_{\text{t}}}$. Then, we calculate the score of user semantic interest embedding and item text embedding $\hat{y_{k}^{\text{t}}}=\text{softmax}({\mathbf{H}_{\text{ex}}^{{}^{\prime}}}^{\top}\mathbf{t}_{k})$, where $\mathbf{t}_{k}\in\mathbb{R}^{d_{\text{t}}}$ is item text embeddings. Then, we use cross-entropy as the loss function for the semantic prediction task.
(9) $\mathcal{L}_{\text{S}}=\sum_{k=1}^{N}{{y_{k}}^{\text{t}}log(\hat{{y_{k}}}^{\text{t}}}),$
3.5.2. Modality Alignment Task
To ensure that the embeddings of different modalities can be in a unified embedding space, we specially designed a modality alignment task. Specifically, the task achieves alignment between behavioral interest representation and semantic interest representation as well as alignment between item ID labels and corresponding text labels through contrastive learning and cosine similarity.
(10) $\text{CL}(\mathbf{e}_{a},\mathbf{e}_{b})=-\frac{1}{N}\sum_{k=1}^{N}log\left(\frac{\text{exp}(\text{sim}(\mathbf{e}_{a}^{k},\mathbf{e}_{b}^{k})/\tau)}{\sum_{j=1}^{N}\text{exp}(\text{sim}(\mathbf{e}_{a}^{k},\mathbf{e}_{b}^{j})/\tau)}\right),$
(11) $\text{COS}(\mathbf{e}_{a},\mathbf{e}_{b})=\frac{1}{N}\sum_{k=1}^{N}\left(1-\text{sim}(\mathbf{e}_{a}^{k},\mathbf{e}_{b}^{k})\right),$
where $\text{sim}(a,b)=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$is the calculation method of cosine similarity, the loss function for the final modality alignment task we denote as follows,
(12) $\mathcal{L}_{\text{A}}=\alpha(\text{CL}(\mathbf{H}_{\text{ex}}^{{}^{\prime}},\mathbf{H}_{\text{im}}^{{}^{\prime}})+\text{CL}(\mathbf{t}_{\text{Tar}},\mathbf{e}_{\text{Tar}}))\\ +\beta(\text{COS}(\mathbf{H}_{\text{ex}}^{{}^{\prime}},\mathbf{H}_{\text{im}}^{{}^{\prime}})+\text{COS}(\mathbf{t}_{\text{Tar}},\mathbf{e}_{\text{Tar}})),$
where $\alpha$ and $\beta$ are hyperparameters that control the ratio of different loss calculation methods. In addition, $12*(\alpha+\beta)=$ 1 means that the sum of the two modal ratio coefficients is 1.